# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v2

@@ Abstract
Large Language Models have demonstrated strong performance across a wide range of tasks, but adapting them efficiently to new domains remains a key challenge. Parameter-Efficient Fine-Tuning (PEFT) methods address this by introducing lightweight, trainable modules while keeping most pre-trained weights fixed. The prevailing approach, LoRA, models updates using a low-rank decomposition, but its expressivity is inherently constrained by the rank. Recent methods like HiRA aim to increase expressivity by incorporating a Hadamard product with the frozen weights, but still rely on the structure of the pre-trained model. We introduce ABBA, a new PEFT architecture that reparameterizes the update as a Hadamard product of two independently learnable low-rank matrices. In contrast to prior work, ABBA fully decouples the update from the pre-trained weights, enabling both components to be optimized freely. This leads to significantly higher expressivity under the same parameter budget. We formally analyze ABBA’s expressive capacity and validate its advantages through matrix reconstruction experiments. Empirically, ABBA achieves state-of-the-art results on arithmetic and commonsense reasoning benchmarks, consistently outperforming existing PEFT methods by a significant margin across multiple models. Our code is publicly available at: https://github.com/CERT-Lab/abba.
@@ 1 Introduction
Large Language Models (LLMs) have become the backbone of modern NLP systems [1, 2, 3, 4, 5, 6, 7], demonstrating strong generalization across a wide range of tasks [8, 9]. However, adapting these models to new tasks typically requires full fine-tuning (FT), which is computationally and memory intensive. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by introducing a small number of trainable parameters while keeping the majority of model weights frozen [10, 11, 12]. Among PEFT approaches, Low-Rank Adaptation (LoRA) [10] is the most widely adopted due to its simplicity and effectiveness. It models the weight update $\Delta W$ as the product of two low-rank matrices, providing a compact parameterization. However, this formulation inherently constrains updates to a low-dimensional subspace, limiting expressivity. Several extensions attempt to overcome this limitation: LoRA-XS [12] restricts updates to a frozen, SVD-derived subspace and training only a small higher-rank intermediary matrix while DoRA [11] modifies only the directional component of the weights via low-rank updates. Despite these architectural variations, all retain LoRA’s core constraint: updates remain strictly low-rank and limited in expressivity.
HiRA [13] addresses LoRA’s limited expressivity by applying a Hadamard product between a low-rank update and the frozen pre-trained weights $W_{0}$, enabling updates that can, in principle, attain full rank. However, HiRA’s expressivity remains tightly coupled to $W_{0}$, as the learned adapters merely modulate the pre-trained weights rather than generating the full update independently. For example, if the target update equals $\mathrm{diag}(W_{0})$, HiRA must learn adapters that approximate the identity matrix, an orthonormal structure that is challenging for low-rank modules to represent accurately (Figure 2).
@@
We support our formulation with both theoretical and empirical evidence. First, we analyze expressivity through a matrix reconstruction task, where Hadamard-structured updates consistently outperform standard LoRA decompositions under the same parameter budget (Figure 2). This demonstrates that ABBA can represent a broader class of updates than LoRA within identical constraints. Next, a toy MNIST experiment (Figure 1) shows that ABBA converges to a solution significantly closer to the true optimum compared to LoRA and HiRA, indicating that its increased expressivity is also practically accessible during learning. We present an exact reformulation of the ABBA update using Khatri–Rao matrix factorization, enabling efficient implementation without approximation. Empirically, ABBA consistently outperforms existing PEFT methods across a broad range of tasks and models, all within the same or lower parameter budget. Our key contributions are summarized as:
@@
  • We provide theoretical and empirical analyses of ABBA’s expressivity, showing that Hadamard-based decomposition consistently outperforms standard low-rank methods in matrix reconstruction.
@@ 2.1 Preliminaries
LoRA [10]. LoRA mitigates this by modeling the update as a low-rank decomposition: $\Delta W=sBA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and $s$ is a scaling factor. This reduces the number of trainable parameters to $r(m+n)$, with $r\ll\min(m,n)$. LoRA can represent any update of rank at most $r$, but cannot express higher-rank updates. Moreover, the projected gradient onto the weight space is also low-rank. While effective for simpler tasks, this limitation becomes significant in settings requiring high-rank updates or gradients [14, 15].
HiRA (Hadamard High-Rank Adaptation)  [13]. HiRA lifts LoRA’s rank limitation by modulating its low-rank update with an element-wise (Hadamard) product with the frozen pre-trained weight $W_{0}$:
@@ 2.2 Improving the Expressivity of HiRA
2.3 ABBA: Highly Expressive and High Rank, Yet Efficient
@@
By the Eckart–Young–Mirsky (EYM) theorem [16, 17], the truncated SVD yields the optimal rank-$r_{1}$ approximation of $W_{0}$. This hybrid initialization anchors the update close to a meaningful low-rank subspace, while enabling the second adapter pair to explore task-specific directions during training. We validate the effectiveness of this strategy empirically in Section 4.1.
@@
Theorem 1 (Khatri–Rao Factorization [18]).
Let $B_{1}A_{1},B_{2}A_{2}\in\mathbb{R}^{m\times n}$ Then, $(B_{1}A_{1})\odot(B_{2}A_{2})=\underbrace{(B_{1}\odot_{r}B_{2})}_{m\times r_{1}r_{2}}\underbrace{(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}}_{r_{1}r_{2}\times n},$ where $\odot_{r}$111 Given $U,V\in\mathbb{R}^{m\times n}$, the row-wise Khatri–Rao product $U\odot_{r}V\in\mathbb{R}^{m\times n^{2}}$ is defined by $[U\odot_{r}V]_{i}:=\left[U_{i1}V_{i},\,U_{i2}V_{i},\,\dots,\,U_{in}V_{i}\right],$ where $V_{i}$ is the $i$-th row of $V$. denotes the row-wise Khatri–Rao product.
@@
Scaling Factor in the ABBA Update.
Theorem 1 reformulates the ABBA update as a Khatri–Rao product, giving it a superficially LoRA-like structure. rsLoRA [19] argues that stable LoRA training requires a scaling factor $\alpha_{\text{LoRA}}\propto 1/\sqrt{r}$. In ABBA, however, the Hadamard structure enables an effective rank of $r_{1}r_{2}$, suggesting an analogous scaling of $1/\sqrt{r_{1}r_{2}}$. Yet, since the parameterizations lie on fundamentally different manifolds, the justification used in rsLoRA does not directly apply. We formally derive the appropriate scaling for ABBA in Theorem 2, and support it with empirical evidence in Section 4.2.
Theorem 2 (Effective Scaling for ABBA Updates).
To maintain training stability and match the update magnitude of LoRA, the ABBA update $\Delta W=s_{\text{ABBA}}\cdot(B_{1}A_{1}\odot B_{2}A_{2})$ should use the scaling factor $s_{\text{ABBA}}=\frac{\alpha_{\text{LoRA}}^{2}}{\sqrt{r_{1}r_{2}}},$ where $\alpha_{\text{LoRA}}$ is the standard LoRA scaling coefficient and $r_{1},r_{2}$ are the ranks of the two ABBA adapter pairs.
Proof.
See Appendix D.2. ∎
2.4 Formalizing and Measuring Expressivity of ABBA
@@
In LoRA, the weight update is modeled as a low-rank decomposition $\Delta W\approx BA$. For any matrix $M\in\mathbb{R}^{m\times n}$, the reconstruction error of this approximation is defined as:
@@
and is lower-bounded by the classical EYM theorem [16, 17], which states that the optimal rank-$r$ approximation is given by the truncated SVD. Since a LoRA adapter of rank $r$ can only represent updates with rank at most $r$, EYM provides a theoretical minimum for $\mathcal{E}(r)$. LoRA achieves this bound exactly when the learned adapters align with the top singular components of $M$; otherwise, practical considerations such as suboptimal initialization may lead to a performance gap.
@@
Previous work [20] establishes a loose upper bound $E_{\text{ABBA},r}\leq E_{\text{LoRA},r}$, which holds trivially by setting one ABBA factor to the rank-$r$ SVD and the other to an all-ones matrix. Empirically, however, they observe a stronger trend: $E_{\text{ABBA},r}\lesssim E_{\text{LoRA},2r},$ which suggests that ABBA can match or outperform a rank-$2r$ SVD approximation using only rank $r$. However, this behavior is not guaranteed for arbitrary matrices, as the quality of reconstruction depends on the spectral properties and structure of the matrix. The only known theoretical comparison between the two is the bound: $E_{\text{LoRA},2r}-E_{\text{ABBA},r}\leq\sum_{i=2r+1}^{r^{2}}\sigma_{i}^{2},$ where $\sigma_{i}$ are the singular values of $M$. While this bound offers some insight, it is loose and does not guarantee a strict ordering between the reconstruction errors.
To better understand practical behavior, we empirically evaluate the reconstruction error of different parameterizations across diverse matrix types. As shown in Figure 2, the ABBA-based Hadamard reparameterization consistently achieves lower reconstruction error than standard LoRA, indicating greater expressivity. This aligns very well with prior work leveraging Hadamard structures for efficient and expressive matrix representations [21], further validating our formulation. Beyond these analyses, we also provide a probabilistic guarantee in Theorem 3 demonstrating when ABBA is likely to outperform LoRA in the presence of structured signals corrupted by noise.
Theorem 3 (Probabilistic Advantage of ABBA over LoRA).
Let $M=M_{\mathrm{true}}+\sigma Z$, where $M_{\mathrm{true}}\in\mathcal{M}_{\mathrm{Had}}$ lies on the Hadamard-product manifold (space in which ABBA updates lie - see Defn. (2)), and $Z\in\mathbb{R}^{m\times n}$ has entries i.i.d with each distributed $\mathcal{N}(0,1)$. Let $\epsilon_{\mathrm{bias}}^{2}:=\|M_{\mathrm{true}}-M_{\mathrm{SVD},r}\|_{F}^{2}$ and $\delta=\exp\left(-\frac{\epsilon_{\mathrm{bias}}^{2}}{8\sigma^{2}}\right)$, then, $\mathbb{P}\left(\|M-M_{\mathrm{ABBA},r}\|_{F}^{2}<\|M-M_{\mathrm{SVD},r}\|_{F}^{2}\right)\geq 1-\delta.$ (7)
Proof.
See Appendix D.3. ∎
@@ 2.5 Stability of the ABBA Update
The scaling factor $s_{\text{ABBA}}$ is critical in controlling the optimization dynamics of the ABBA update. It must be chosen carefully to scale appropriately with the ranks $r_{1}$ and $r_{2}$ of the underlying low-rank factors. If set too low, learning stagnates; if too high, training may diverge. While ABBA resembles LoRA in structure, its effective rank is $r_{1}r_{2}$, and the scaling behavior must reflect this increased capacity. Inspired by the scaling analysis in rsLoRA [19] and LoRA-GA [22], where stability is achieved under a complexity of $\mathcal{O}(1)$ with respect to rank, one might expect similar scaling for ABBA. However, since ABBA and LoRA inhabit different parameter spaces, these arguments do not transfer directly. To formalize this, we introduce the notion of rank-stability for ABBA in Definition 1, which ensures that forward and backward dynamics remain well-conditioned as $r_{1}$ and $r_{2}$ vary.
Definition 1 (Rank Stability of ABBA Adapters [19, 22]).
@@
Theorem 4 (Rank-Stability of ABBA Initialization).
Let $s_{\text{ABBA}}$ be the scaling factor from Theorem 2, and let the initialization follow Section 2.3. Then the ABBA update satisfies the forward and backward stability conditions of Definition 1, and is thus rank-stable.
@@ 3 Experiments
We evaluate ABBA on a range of models, specifically Llama-3.2 1B [23], Llama-3.2 3B [23], Mistral-7B [24], and Gemma-2 9B [25], to test its effectiveness across diverse scales and architectures. We run experiments on multiple benchmarks to capture varied trends. Appendix H details our training configurations, and Appendix I lists dataset specifics. For fair comparison with LoRA and other PEFT methods, we match the number of trainable parameters in ABBA by setting $r_{1}=r_{2}=r/2$.
Baselines. We compare ABBA against full fine-tuning, LoRA [10], and several strong LoRA variants: rsLoRA [19], PiSSA [26], DoRA [11], LoRA-Pro [14], and HiRA [13].
@@ 3.1 Commonsense Reasoning
We fine-tune LLaMA-3.2 models at 1B and 3B scales [23] on CommonSense170K, a multi-task dataset comprising eight commonsense reasoning benchmarks [27]. These include OBQA [28], ARC-Challenge and ARC-Easy [29], WinoGrande [30], HellaSwag [31], PIQA [32], SIQA [33], and BoolQ [34]. We evaluate performance on each dataset independently to capture task-specific generalization. We insert LoRA modules into the key, query, and value projections, the attention output, and all feedforward layers. Table 1 reports the results. ABBA consistently outperforms all other PEFT methods across both models, and in many cases surpasses full FT.

### v3

# ABBA-Adapters: Efficient and Expressive Fine-Tuning of Foundation Models

@@ Abstract
Large Language Models have demonstrated strong performance across a wide range of tasks, but adapting them efficiently to new domains remains a key challenge. Parameter-Efficient Fine-Tuning (PEFT) methods address this by introducing lightweight, trainable modules while keeping most pre-trained weights fixed. The prevailing approach, LoRA, models updates using a low-rank decomposition, but its expressivity is inherently constrained by the rank. Recent methods like HiRA aim to increase expressivity by incorporating a Hadamard product with the frozen weights, but still rely on the structure of the pre-trained model. We introduce ABBA, a new PEFT architecture that reparameterizes the update as a Hadamard product of two independently learnable low-rank matrices. In contrast to prior work, ABBA fully decouples the update from the pre-trained weights, enabling both components to be optimized freely. This leads to significantly higher expressivity under the same parameter budget, a property we validate through matrix reconstruction experiments. Empirically, ABBA achieves state-of-the-art results on arithmetic and commonsense reasoning benchmarks, consistently outperforming existing PEFT methods by a significant margin across multiple models. Our code is publicly available at: https://github.com/CERT-Lab/abba.
@@ 1 Introduction
Large Language Models (LLMs) have become the backbone of modern NLP systems (1, 2, 3, 4, 5, 6, 7), demonstrating strong generalization across a wide range of tasks (8, 9). However, adapting these models to new tasks typically requires full fine-tuning (FT), which is computationally and memory intensive. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by introducing a small number of trainable parameters while keeping the majority of model weights frozen (10, 11, 12). Among PEFT approaches, Low-Rank Adaptation (LoRA) (10) is the most widely adopted due to its simplicity and effectiveness. It models the weight update $\Delta W$ as the product of two low-rank matrices, providing a compact parameterization. However, this formulation inherently constrains updates to a low-dimensional subspace, limiting expressivity. Several extensions attempt to overcome this limitation: LoRA-XS (12) restricts updates to a frozen, SVD-derived subspace and training only a small higher-rank intermediary matrix while DoRA (11) modifies only the directional component of the weights via low-rank updates. Despite these architectural variations, all retain LoRA’s core constraint: updates remain strictly low-rank and limited in expressivity.
HiRA (13) addresses LoRA’s limited expressivity by applying a Hadamard product between a low-rank update and the frozen pre-trained weights $W_{0}$, enabling updates that can, in principle, attain full rank. However, HiRA’s expressivity remains tightly coupled to $W_{0}$, as the learned adapters merely modulate the pre-trained weights rather than generating the full update independently. For example, if the target update equals $\mathrm{diag}(W_{0})$, HiRA must learn adapters that approximate the identity matrix, an orthonormal structure that is challenging for low-rank modules to represent accurately (Figure 2).
@@
We analyze expressivity through a matrix reconstruction task, where Hadamard-structured updates consistently outperform standard LoRA decompositions under the same parameter budget (Figure 2). This demonstrates that ABBA can represent a broader class of updates than LoRA within identical constraints. A toy MNIST experiment (Figure 1) shows that ABBA converges to a solution significantly closer to the true optimum compared to LoRA and HiRA, indicating that its increased expressivity is also practically accessible during learning. We present an exact reformulation of the ABBA update using the Khatri–Rao matrix factorization, enabling efficient implementation without approximation. Empirically, ABBA consistently outperforms existing PEFT methods across a broad range of tasks and models, all within the same or lower parameter budget. In addition, we conduct extensive ablation studies to validate the effectiveness of our design choices and hyperparameter settings. Our key contributions are summarized as:
@@
  • We provide empirical analyses of ABBA’s expressivity, showing that Hadamard-based decomposition consistently outperforms standard low-rank methods in matrix reconstruction.
@@ 2.1 Preliminaries
LoRA (10). LoRA mitigates this by modeling the update as a low-rank decomposition: $\Delta W=sBA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, and $s$ is a scaling factor. This reduces the number of trainable parameters to $r(m+n)$, with $r\ll\min(m,n)$. LoRA can represent any update of rank at most $r$, but cannot express higher-rank updates. Moreover, the projected gradient onto the weight space is also low-rank. While effective for simpler tasks, this limitation becomes significant in settings requiring high-rank updates or gradients (14, 15).
HiRA (Hadamard High-Rank Adaptation)  (13). HiRA lifts LoRA’s rank limitation by modulating its low-rank update with an element-wise (Hadamard) product with the frozen pre-trained weight $W_{0}$:
@@ 2.2 Improving the Expressivity of HiRA
2.3 ABBA-Adapters: Improved Expressivity and High Rank, Yet Efficient
@@
By the Eckart–Young–Mirsky (EYM) theorem (16, 17), the truncated SVD yields the optimal rank-$r_{1}$ approximation of $W_{0}$. This hybrid initialization anchors the update close to a meaningful low-rank subspace, while enabling the second adapter pair to explore task-specific directions during training. We validate the effectiveness of this strategy empirically in Section 4.1.
@@
Theorem 1 (Khatri–Rao Factorization (18)).
Let $B_{1}A_{1},B_{2}A_{2}\in\mathbb{R}^{m\times n}$ . Then, $(B_{1}A_{1})\odot(B_{2}A_{2})=\underbrace{(B_{1}\odot_{r}B_{2})}_{m\times r_{1}r_{2}}\underbrace{(A_{1}^{\top}\odot_{r}A_{2}^{\top})^{\top}}_{r_{1}r_{2}\times n},$ where $\odot_{r}$ [^1] Given $U,V\in\mathbb{R}^{m\times n}$, the row-wise Khatri–Rao product $U\odot_{r}V\in\mathbb{R}^{m\times n^{2}}$ is defined by $U\odot_{r}V]_{i}:=\left[U_{i1}V_{i},\,U_{i2}V_{i},\,\dots,\,U_{in}V_{i}\right],$ where $V_{i}$ is the $i$-th row of $V$. denotes the row-wise Khatri–Rao product.
@@
2.4 Expressivity of ABBA
@@
In LoRA, the weight update is modeled as a low-rank decomposition $\Delta W=BA$. For any matrix $M\in\mathbb{R}^{m\times n}$, the reconstruction error of this approximation is defined as:
@@
and is lower-bounded by the classical EYM theorem (16, 17), which states that the optimal rank-$r$ approximation is given by the truncated SVD. Since a LoRA adapter of rank $r$ can only represent updates with rank at most $r$, EYM provides a theoretical minimum for $\mathcal{E}(r)$. LoRA achieves this bound exactly when the learned adapters align with the top singular components of $M$; otherwise, practical considerations such as suboptimal initialization may lead to a performance gap.
@@
Prior work (19) establishes a loose upper bound $E_{\text{ABBA},r}\leq E_{\text{LoRA},r}$, which holds trivially by setting one ABBA factor to the rank-$r$ SVD and the other to an all-ones matrix. Empirically, however, they observe a stronger trend: $E_{\text{ABBA},r}\lesssim E_{\text{LoRA},2r},$ which suggests that ABBA can match or outperform a rank-$2r$ SVD approximation using only rank $r$. However, this is not guaranteed for arbitrary matrices, as the quality of reconstruction depends on the spectral properties and structure of the matrix. The only known theoretical comparison between the two is the bound: $E_{\text{LoRA},2r}-E_{\text{ABBA},r}\leq\sum_{i=2r+1}^{r^{2}}\sigma_{i}^{2},$ where $\sigma_{i}$ are the singular values of $M$ (19). While this bound offers some insight, it is loose and does not guarantee a strict ordering between the reconstruction errors.
To better understand practical behavior, we empirically evaluate the reconstruction error of different parameterizations across diverse matrix types. As shown in Figure 2, the ABBA-based Hadamard reparameterization consistently achieves lower reconstruction error than standard LoRA, indicating greater expressivity. This aligns very well with prior work leveraging Hadamard structures for efficient and expressive matrix representations (20), further validating our formulation.
@@ 2.5 Stability of the ABBA Update
The scaling factor $s$ is critical in controlling the optimization dynamics of the ABBA update. It must be chosen carefully to scale appropriately with the ranks $r_{1}$ and $r_{2}$ of the underlying low-rank factors. If set too low, learning stagnates; if too high, training may diverge. While ABBA resembles LoRA in structure, its effective rank is $r_{1}r_{2}$, and the scaling behavior must reflect this increased capacity. Inspired by the scaling analysis in rsLoRA (21), where stability is achieved under a complexity of $\mathcal{O}(1)$ with respect to rank, one might expect similar scaling for ABBA. However, since ABBA and LoRA inhabit different parameter spaces, these arguments do not transfer directly. To formalize this, we introduce the notion of rank-stability for ABBA in Definition 1, which ensures that forward and backward dynamics remain well-conditioned as $r_{1}$ and $r_{2}$ vary.
Definition 1 (Rank Stability of ABBA Adapters (21, 22)).
@@
Theorem 2 (Rank-Stability of ABBA).
Let the ABBA update be $\Delta W=s_{\text{ABBA}}\,(B_{1}A_{1})\odot(B_{2}A_{2}),$ with $B_{1},A_{1},B_{2},A_{2}$ independent, mean-zero random matrices with finite variance. Then, the update satisfies the stability conditions of Definition 1 if and only if: $s_{\text{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right).$
@@
Final ABBA Parameterization.
We express the ABBA scaling factor as $s_{\mathrm{ABBA}}\;=\;\frac{\alpha^{2}}{\sqrt{r_{1}r_{2}}},$ since Theorem 2 shows that $s_{\mathrm{ABBA}}\in\Theta\!\left(\tfrac{1}{\sqrt{r_{1}r_{2}}}\right)$. With this choice, the ABBA update is:
$\Delta W\;=\;\frac{\alpha^{2}}{\sqrt{r_{1}r_{2}}}\,(B_{1}A_{1})\odot(B_{2}A_{2}).$ (7)
Equivalently, to highlight the connection with LoRA, ABBA can be expressed as the Hadamard product of two rank-stabilized LoRA adapters:
$\Delta W\;=\;\left(\frac{\alpha}{\sqrt{r_{1}}}B_{1}A_{1}\right)\odot\left(\frac{\alpha}{\sqrt{r_{2}}}B_{2}A_{2}\right).$ (8)
@@ 3 Experiments
We evaluate ABBA on a range of models, specifically Llama-3.2 1B (23), Llama-3.2 3B (23), Mistral-7B (24), and Gemma-2 9B (25), to test its effectiveness across diverse scales and architectures. We run experiments on multiple benchmarks to capture varied trends. Appendix H details our training configurations, and Appendix I lists dataset specifics. For fair comparison with LoRA and other PEFT methods, we match the number of trainable parameters in ABBA by setting $r_{1}=r_{2}=r/2$.
Baselines. We compare ABBA against full fine-tuning, LoRA (10), and several strong LoRA variants: rsLoRA (21), PiSSA (26), DoRA (11), LoRA-Pro (14), and HiRA (13).
@@ 3.1 Commonsense Reasoning
We fine-tune Llama-3.2 models at 1B and 3B scales (23) on CommonSense170K, a multi-task dataset comprising eight commonsense reasoning benchmarks (27). These include OBQA (28), ARC-Challenge and ARC-Easy (29), WinoGrande (30), HellaSwag (31), PIQA (32), SIQA (33), and BoolQ (34). We evaluate performance on each dataset independently to capture task-specific generalization. We insert LoRA modules into the key, query, and value projections, the attention output, and all feedforward layers. Table 1 reports the results. ABBA consistently outperforms all other PEFT methods across both models, and in many cases surpasses full FT.